OpenStudy (anonymous):

**********HELP********** What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)?

OpenStudy (lncognlto):

First you need to find the gradient, m = (y2-y1)/(x2-x1) Then subst m and a point into y = mx + c. Solve for c, then subst. c and m in y = mx +c. This would be the answer.